Operations Report – January 6th

Crew 272 Operations Report 6-1-2023

SOL: 5

Name of person filing report: Mason Kuhn

Non-nominal systems: Headset in box 3

Notes on non-nominal systems: Radio 3 has been tested with other headsets and is nominal. Headset 3 has poor audio and others can only hear static from headset 3.


Spirit rover used: Yes

Hours: 206.8

Beginning charge: 100%

Ending charge: 71%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 111.2

Beginning charge: 99%

Ending charge: 81%

Currently charging: Yes

Curiosity rover used: No

Hours: 218.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: Spirit and Opportunity were used on the EVA. It was noticed that Opportunity held a charge much better than Spirit did, especially in the early stages of the EVA.

Summary of Hab operations:

WATER USE: 13.1 gallons

Water (static tank): 298.4 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits 7, 8, 10, 11 were used on today’s EVA. All suits are currently charging. Suit 7 was charged to 13.68 V before the EVA, but the crew member wearing it noted that the airflow was poor. I will look to troubleshoot suit 7 tonight. All radios performed nominally during the EVA. The headset in box 7 was occasionally faulty during the EVA, but it could have been an issue of not pressing the comms button with enough force. Headset 3 was tested prior to EVA and was not functional (described in non-nominal systems).

Summary of GreenHab operations: Watered and monitored all GreenHab crops. Watered experimental crops new crops in regolith & soil experiments. Planted lettuce and kale. Harvested beans, chives, and microgreens.

WATER USE: 9 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 11g beans, 3g chives, 10g microgreens

Summary of ScienceDome operations:

Dual split: Will run overnight (approximately 8 pm – 8 am)

Summary of RAM operations: The CO alarm’s low battery chirp did not return during the day or night time. No items were retrieved from the RAM.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support:

  • Request for information on if the crew can have items delivered to MDRS
    • Specifically, a USB-C to lightning cable (Skydio brand only)
  • Request for new salt-shaker & new dishwashing sponge
Copyright © The Mars Society. All rights reserved. | Main Site